Bexley:

the diameter of circle j is 18 cm. the diameter of circle k is 31 cm. Describe how you can compare the areas of the two circles

Renne:

To compare the areas of circle j and circle k, you can use the formula for the area of a circle, which is A = πr^2. Since we know the diameter of each circle, we can find the radius by dividing the diameter by 2. For circle j, the radius is 9 cm (18 cm / 2), and for circle k, the radius is 15.5 cm (31 cm / 2). Next, we can plug the radius of each circle into the formula to find their respective areas. For circle j, the area is A = π(9 cm)^2 ≈ 254.47 cm^2. For circle k, the area is A = π(15.5 cm)^2 ≈ 755.14 cm^2. Therefore, we can see that the area of circle k is larger than the area of circle j.

Renne:

@tbone @arieonna am I correct?

surjithayer:

for J \[area~A_{1}=\pi(\frac{ 18 }{ 2 })^2=\frac{ 324\pi }{ 4 }\] for k \[area~A _{2}={ \pi(\frac{ 31 }{ 2 })^2 }=\frac{961\pi}{4}\] divide \[\frac{ A _{2} }{ A _{1}}=\frac{ \frac{ 961\pi }{ 4 } }{ \frac{ 324\pi }{ 4 } }=\frac{ 961 }{ 324 }\]

Tbone:

@renne wrote:
@tbone @arieonna am I correct?
TvT STILL AINT LEARNED DIAMETER

Renne:

@tbone wrote:
@renne wrote:
@tbone @arieonna am I correct?
TvT STILL AINT LEARNED DIAMETER
lmaooo I'm dyinggg

Bexley:

@renne wrote:
To compare the areas of circle j and circle k, you can use the formula for the area of a circle, which is A = πr^2. Since we know the diameter of each circle, we can find the radius by dividing the diameter by 2. For circle j, the radius is 9 cm (18 cm / 2), and for circle k, the radius is 15.5 cm (31 cm / 2). Next, we can plug the radius of each circle into the formula to find their respective areas. For circle j, the area is A = π(9 cm)^2 ≈ 254.47 cm^2. For circle k, the area is A = π(15.5 cm)^2 ≈ 755.14 cm^2. Therefore, we can see that the area of circle k is larger than the area of circle j.
Thank you that is really helpful :)